Shift Summary for owl shift, 09.14.99

Jim Graham (jagraham@hep.uchicago.edu)
Tue, 14 Sep 1999 07:04:03 -0500

Shift Summary for owl shift, 09.14.99

On Shift: JAG, MA

Hours of beam delivered at all intensities: 6.5 ( / 7.0)
Hours of beam delivered at nominal intensity: 6.5
Hours of E832 data (special runs not included): 6.0
Hours of special runs (laser scans, muon runs...): 0.1

Accounting of lost time: Beam vanishes, tape changes and magnet ramping,
kfrends.

========================================================================
Runs ended this shift:
Run 14516 - 23 M events - BEAM_832
Run 14517 - 350 k events - BEAM_MUON

Tapes brought to FCC this shift:

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Good 832 running, with some aborts. Did muon run and magnet polarity
flip in the middle.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14516 320 00:00 inherit from eve shift
01:12 Trigger hangs on busy from AZLATs; kfrend
01:18 mount onsplit tape QKCC20
01:19 beam gone for TeV station water heater
01:37 beam back -- they lost a TeV RF unit.
lots of aborts in here

471 03:58 Call from MCR -- they will try to fix the
large number of aborts. BOA will probably increase.
BOA up to 1.6
04:10 BOA chilled out to 1.3 -- 1.4
554 05:52 tapes full, run ends. Update ADGN constants as per
Julie's request. Change tapes

14517 06:02 Go to muon mode. Do drt_fresh_start to restart DAQ
monitoring tool.
06:13 End muon run

14518 06:25 Start E832 running again.

========================================================================
Follow up Notes on Previous Shift

Updated ADGN constants, did muon run and polarity flip.

========================================================================
Items to Follow up on Next Shift

Remember that reg HV is no longer interlocked to vacuum. If vacuum
fails, one must manually turn off reg HV. See note next to alarm screen
terminal.

Regards, Jim ( for Mike)